# MEGA-Bench[Uncaptioned image]: Scaling Multimodal Evaluation to over 500 Real-World Tasks

### v2

@@ 1 Introduction
Large foundation models (OpenAI, 2023; 2024a; Anthropic, 2024a; Google, 2023; Meta, 2024; Alibaba, 2024) have dramatically transformed the landscape of artificial intelligence by showcasing exceptional capabilities across various tasks and domains. Originating in the realm of natural language processing, these models have progressively expanded to perceive and interpret multimodal information, including single images, multiple images, and videos. Previously, multimodal models were mainly used for standardized tasks like image captioning (Lin et al., 2014), video captioning (Wang et al., 2019), and visual question answering (Antol et al., 2015; Goyal et al., 2017; Xiao et al., 2021). With the recent progress on multimodal alignment, these models have shown great potential to solve any desired task with a well-designed prompt. As a result, people have applied them to assist with many realistic tasks like “web navigation” (Koh et al., 2024), “game playing” (Valevski et al., 2024), “travel planning” (Xie et al., 2024), “visual navigation” (Wang et al., 2023a), “sports analysis” (Xia et al., 2024), “visual entity recognition” (Hu et al., 2023), “visual quality assessment” (Ku et al., 2024), and more. These efforts have significantly increased the utility of multimodal models.
@@
- Expensive Inference Cost: The full evaluation process is expensive in the computation cost, the time elapsed, and API price in \$. Since many examples or tasks are highly similar in the capabilities that they assess, overly repetitive evaluation at a large scale causes a waste of resources.
- Unmanageable Setups: Each benchmark has its own complexities when setting up the evaluation. For example, VQA (Goyal et al., 2017) has four splits including val, dev-test, std-test, and test. It is hard to track the exact setup of different baseline models to ensure a fair comparison.
@@
Based on the MEGA-Bench, comprehensive studies on existing flagship and efficiency multimodal foundation models (covering most open-source software and proprietary model APIs) have discovered the following findings:
1. Among flagship models, Claude 3.5 Sonnet (1022) and GPT-4o (0513) currently lead in performance across a wide range of multimodal tasks, with less than a 0.1% difference in their overall scores. Detailed breakdown analyses show Claude 3.5 Sonnet excels in planning and math, while GPT-4o leads in information extraction and knowledge-intensive tasks.
@@
4. Proprietary models can effectively leverage Chain-of-Thought (CoT) prompting to improve their performance, while open-source models hardly produce helpful reasoning processes. In our evaluation results, 11 out of 16 open-source models get worse results with CoT prompting.
@@ 2 Related Work
Multimodal benchmarks. Benchmarking in vision-language models has been a long-standing research problem. Prior to large multimodal models, most benchmarks were focused on specific tasks or skills. Some benchmarks like VQA (Antol et al., 2015), GQA (Hudson & Manning, 2019), and ViswizVQA (Gurari et al., 2018) focus on photograph or natural images. ChartQA (Masry et al., 2022), InfoVQA (Mathew et al., 2022), DocVQA (Mathew et al., 2021), and OCR-VQA (Mishra et al., 2019) focus more on documents, infographics, and other similar media. Later on, there was a trend to build more well-rounded benchmarks to cover wider range of skills including MMMU (Yue et al., 2024a), MMMU-Pro (Yue et al., 2024b), MMT-Bench (Ying et al., 2024), MMBench (Liu et al., 2023b), ScienceQA (Lu et al., 2022), and more. However, due to the diversity of these different tasks, most benchmarks simply use multiple-choice questions for all of the problems. Therefore, these benchmarks are not able to reflect the generational abilities of multimodal models. Complementary to this, LMsys arena (Chiang et al., 2024) and WildVision arena (Lu et al., 2024) have proposed to use user voting and Elo-ranking to benchmark multimodal models. Our benchmark is the first to scale up the tasks by a signficant magnitude. Also, our benchmark provides a breakdown report to show the models’ abilities across multiple dimensions.
@@ 3 MEGA-Bench
MEGA-Bench is a comprehensive multimodal benchmark that spans 7 input formats, 6 output formats, 10 different types of skills, and a varying number of visual inputs, whether single-image, multi-image, or from video, as shown in Figure 2. Our benchmark covers 8 distinct subject areas in a hierarchical taxonomy to evaluate the ability of VLMs to tackle a variety of different tasks.
@@ 3.1 Benchmark Construction Process
Preparation. Figure 3 illustrates our annotation process. In the conceptualization stage, we propose a “draft” task taxonomy tree with the top two levels of Figure 1. The first level consists of the task scopes like “Perception”, “Planning”, “Reasoning”, etc. The second level consists of more concrete categories like “Document Understanding”, “App Understanding”, “Logic Reasoning”, and so forth. We add a few exemplars under each second-level node and write detailed descriptions of the tasks we expect from the annotators.
@@
Figure 3: The annotation process of MEGA-Bench. We first propose a “draft” taxonomy tree, and then distribute the meta-nodes to different annotators. We allow the annotators to gradually refine the tree structure as they add new tasks. Each task consists of many examples and has a shared task-level instruction, several per-example questions, and several per-question ground truth answers.
@@
We then distribute the second-level nodes in “draft” taxonomy tree to the annotators based on their preferences. This top-down annotation framework can help minimize the overlaps of tasks contributed by different annotators. To ensure consistent commitment throughout the project, we call up over 16 designated expert annotators. Our annotators are all graduate students or above from different majors like Computer Science, Electronics, Bio-statistics, and Finance. The annotators can 1) refine the “draft” taxonomy by adding/deleting nodes, 2) add “task group” nodes and then add a series of tasks under that, and 3) directly add tasks under the second-level meta-task node. To simplify the task contribution, we develop tools to facilitate the annotation process, including 1) an interactive annotation tool that defines the annotation format and automatically unifies all annotations as JSON files (Please see §B.1 for our annotation format), 2) a GitHub repository to coordinate the task submission, reviewing, and discussion process, which was inspired by BIG-bench (Srivastava et al., 2022), and 3) a visualization tool that allows annotators to browse the existing tasks and the evaluation results of representative vision language models (VLMs). We coordinate all the annotators to ensure they understand our expectations and continuously improve our tools.
@@
Task annotation. The annotation process contains two rounds. The annotators submit tasks to the benchmark by creating pull requests (PR) to the main branch of our GitHub repository.
In the first round of the annotation process, we ask the annotators to contribute 20 tasks following the principles below to ensure the quality of the task:
@@
$\bullet$ Data source and answer format: Creative tasks with diverse data sources and answer formats are encouraged. If the data was collected from existing datasets, we ask annotators to adapt the original annotation into more specific questions and design more diverse answer formats.
@@
Before the second round of annotations, the core contributors gathered with the annotators to review all the contributed tasks in the taxonomy tree and discover the biases in the task distribution. We then host another brainstorming session to propose new meta-tasks to balance the distribution and maximize the coverage. We then distribute the new meta-tasks to our annotators in the second round and follow the same guidelines as before to finish the rest of the annotations. After this round, each annotator contributes at least another 10 tasks to meet the requirements (total $\geq$ 30).
Quality control and refinement. We leverage state-of-the-art VLMs to help the annotators better examine the quality of the tasks. Concretely, we gather the results of GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro and compute an average score on each task. Tasks with almost 1.0 scores often have trivial questions (based on manual inspection) and can hardly distinguish the ability of different models. We ask the corresponding annotators to investigate and augment those tasks to increase the difficulty. For tasks with almost zero scores, the task reviewers audit them carefully and remove them if the zero score comes from an incorrect annotation or insufficient question context. Finally, the benchmark contains a total 505 tasks with roughly 8,200 examples, which is large enough to minimize the sample variance within each high-level taxonomy node. Please refer to §4.3 for an in-depth analysis of the number of examples per task.
@@ 3.2 Metrics for answers in diverse output formats
Figure 4: Representative examples for MEGA-Bench’s diverse output formats and the corresponding customized metrics. The outputs are the real responses from GPT-4o (OpenAI, 2024a). We implement robust parsing to extract the final answer from raw responses.
To properly evaluate the tasks with different output formats, we develop a set of highly-customized evaluation metrics in parallel with the benchmark construction process (§3.1). Figure 4 shows several examples of the model outputs along with the task’s associated metrics. When new tasks are submitted to our GitHub repository, we implement any new metrics specified by the task authors. We use two types of metrics: rule-based metrics and LLM-assisted metrics.
Rule-based metrics. When there is a unique answer under the question context or the correctness of the answer can be verified by rules (e.g., if the generated story/poetry meets the desired formats, or if the generated code can pass the given test cases), we implement rule-based metrics for evaluation. To satisfy the needs of all tasks submitted by our annotators, we end up with a suite of over 40 rule-based metrics. Robust string parsing is also implemented to extract the answer from the model’s response. We conduct a sanity check to ensure the correct implementation of all rule-based metrics. Specifically, we create an “oracle” model that always returns the ground truth annotation. We then compute its score over all tasks evaluated with rule-based metrics. The sanity check is passed when the “oracle” model gets a 1.0 score. See §D.4 for details.
@@
We divide the tasks into two subsets based on the different evaluation processes. The Core Set is evaluated with rule-based metrics to make the evaluation fast and cost-free. The Open-Ended Set is evaluated with metrics that use an LLM-as-a-judge, where the evaluation pipeline calls a proprietary LLM over an API. Specifically, we use GPT-4o-0806 (OpenAI, 2024a) as the judge LLM. The Core and Open-Ended sets contain 440 and 65 tasks, respectively.
@@ 3.3 Multi-dimensional Keywords for fine-grained analysis
Existing multi-task multimodal benchmarks analyze models according to dimensions like the image type and academic discipline (Yue et al., 2024a), ability (Liu et al., 2023b), or meta-task (Ying et al., 2024). MEGA-Bench offers a broad and diverse range of coverage across all these dimensions, and extends even further beyond them.
As explained in §3.1, the taxonomy tree divides the tasks into general application scenarios, the most manageable dimension for distributing the annotation efforts to different annotators. After we collected all tasks and finished the quality control process, we grouped all tasks based on four extra dimensions: input visual type, input visual number, output format, and required skills (Figure 2). Each dimension has 6 to 10 keywords, enabling fine-grained analysis and comparison. Interactive visualization tools can then be developed based on our evaluation results, which allows model developers to delve deep into different aspects of a model, and compare different models comprehensively like how a scout compares athletes.
@@
GQA (Hudson & Manning, 2019) New Photo 1 Image Phrase/Bool/Num 1 1
@@
MathVista (Lu et al., 2023) Repurposed Existing 1 Image Free-form (FF) 1 31
MMBench (Liu et al., 2023b) Repurposed Existing 1 Image Multi-choice (MC) 1 20
@@
MMVet (Yu et al., 2024b) Repurposed Existing 1 Image Multi-choice (MC) 1 16
MMMU (Yue et al., 2024a) New Diverse $\geq$ 1 Image Multi-choice (MC) 1 30
@@
MileBench (Song et al., 2024) Repurposed Existing $>$ 1 Image MC or FF 2 28
@@
MMTBench (Ying et al., 2024) Repurposed Existing Image/Videos Multi-choice (MC) 1 162
MEGA-Bench New Unrestricted Image/Videos Unrestricted 45 505
@@ 3.4 Dataset statistics and comparison with other benchmarks
MEGA-Bench contains 505 real-world tasks with 8,186 manually annotated or repurposed samples. Even for repurposed data, considerable effort is needed to convert the original annotations into specific task descriptions, diverse output formats, and additional instructions to include auxiliary information and about the formatting. Figure 2 shows the task distribution of all five dimensions, and the detailed breakdown statistics are in §C.2.
Table 1 compares MEGA-Bench to existing multimodal benchmarks. The key feature of our benchmark is the diversity across all aspects in the table. MMMU (Yue et al., 2024a; b) focuses on college-level exam questions with various discipline and image formats, testing visual and textual fusion with advanced knowledge. All questions are single-image and answered in multiple-choice format. MMT-Bench (Ying et al., 2024) covers 162 concrete sub-tasks, enabling in-depth analysis based on their “taskonomy” and diverse input forms. However, all of the tasks MMT-Bench are from existing datasets and these tasks are mostly to under the “Perception” sub-tree in our taxonomy, and all outputs are in multiple-choice form like MMMU. To maximize task coverage and the diversity in model outputs with cost-effective inference, MEGA-Bench includes a much broader range of task types and output formats, while having fewer total samples compared to existing benchmarks.

### v3

@@ 1 Introduction
Large foundation models (OpenAI, 2023; 2024a; Anthropic, 2024a; Google, 2023; Meta, 2024; Alibaba, 2024) have dramatically transformed the landscape of artificial intelligence by showcasing exceptional capabilities across various tasks and domains. Originating in the realm of natural language processing, these models have progressively expanded to perceive and interpret multimodal information, including single images, multiple images, and videos. Previously, multimodal models were mainly used for standardized tasks like image captioning (Lin et al., 2014), video captioning (Wang et al., 2019), and visual question answering (Antol et al., 2015; Goyal et al., 2017; Xiao et al., 2021). With the recent progress on multimodal alignment, these models have shown great potential to solve many diverse and complex tasks with well-designed prompts. As a result, people have applied them to assist with many realistic tasks like “web navigation” (Koh et al., 2024), “game playing” (Valevski et al., 2024), “travel planning” (Xie et al., 2024), “visual navigation” (Wang et al., 2023a), “sports analysis” (Xia et al., 2024), “visual entity recognition” (Hu et al., 2023), “visual quality assessment” (Ku et al., 2024), and more. These efforts have significantly increased the utility of multimodal models.
@@
- Expensive Inference Cost: The full evaluation process is expensive regarding computation cost/time or API expense. Since many examples or tasks are similar in the capabilities they assess (e.g., DocVQA Mathew et al. (2021) alone has thousands of examples for examining doc understanding and OCR-related abilities), overly repetitive evaluation at a large scale leads to resource waste.
- Unmanageable Setups: Each benchmark has complexities when setting up the evaluation. For example, VQA (Goyal et al., 2017) alone has four splits (val, dev-test, std-test, and test). It is hard to track the exact setup of different baseline models to ensure a fair comparison.
@@
Using MEGA-Bench, we conducted comprehensive studies of popular flagship and efficiency models (with both open-source software and proprietary APIs) and identified some findings below:
1. Among flagship models, Claude 3.5 Sonnet (1022) and GPT-4o (0513) currently lead in performance across a wide range of multimodal tasks, with less than a 0.1% difference in their overall scores. Our detailed breakdown shows that Claude 3.5 Sonnet excels in planning and math with its latest upgrade bringing clear boosts in processing UI/Infographics inputs, while GPT-4o leads in information extraction and knowledge-intensive tasks.
@@
4. Proprietary models can effectively leverage Chain-of-Thought (CoT) prompting to improve their performance, while open-source models hardly produce helpful reasoning processes. In our evaluation results, 10 of 13 open-source models get worse results with CoT prompting.
@@ 2 Related Work
Multimodal benchmarks. Benchmarking in vision-language models has been a long-standing research problem. Before the era of large multimodal models, most benchmarks were designed for specific tasks or skills. Some benchmarks like VQA (Antol et al., 2015), GQA (Hudson & Manning, 2019), and ViswizVQA (Gurari et al., 2018) focus on photograph or natural images. ChartQA (Masry et al., 2022), InfoVQA (Mathew et al., 2022), DocVQA (Mathew et al., 2021), and OCR-VQA (Mishra et al., 2019) focus more on documents, infographics, and other similar media. Later on, there was a trend to build more well-rounded benchmarks to cover a wider range of skills or topics, such as ScienceQA (Lu et al., 2022), MMBench (Liu et al., 2023b), MMMU (Yue et al., 2024a; b), MMT-Bench (Ying et al., 2024), and more. However, due to the diversity of these different tasks, most benchmarks use multiple-choice questions for all problems. Therefore, these benchmarks cannot fully reflect the generational abilities of multimodal models. Complementary to this, LMsys arena (Chiang et al., 2024) and WildVision arena (Lu et al., 2024) have proposed to use user voting and Elo-ranking to benchmark multimodal models. Our benchmark is the first to scale up the tasks by a significant magnitude. Furthermore, our benchmark provides a breakdown report to analyze multimodal models across multiple dimensions.
@@ 3 MEGA-Bench
MEGA-Bench is a comprehensive multimodal benchmark that spans 7 input formats, 6 output formats, 10 different types of skills, and a varying number of visual inputs, whether single-image, multi-image or from video, as shown in Figure 2. Our benchmark covers 8 distinct subject areas in a hierarchical taxonomy to evaluate VLMs’ ability to tackle various tasks.
@@ 3.1 Benchmark Construction Process
Preparation. Figure 3 illustrates our annotation process. In the conceptualization stage, we propose a “draft” task taxonomy tree with the top two levels of Figure 1 by getting inspirations from existing multi-task or multi-discipline LLM/VLM benchmarks (Srivastava et al., 2022; Liu et al., 2023b; Yue et al., 2024a). The first level consists of general applications like “perception”, “planning”, “reasoning”, etc., while the second level has more concrete meta-tasks like “document understanding”, “app function understanding”, “logic reasoning”, etc. We host a brainstorming session to add exemplars under each second-level node and write descriptions about the number and quality of the tasks we expect. Based on our empirical observations of how general users use VLMs in real-world scenarios, we assign more task budgets to perception, knowledge, and information extraction than other first-level nodes while strictly monitoring the application-level distribution balance in the annotation process. We then distribute the second-level nodes in the “draft” tree to the annotators. This top-down framework minimizes overlaps between annotators and facilitates overall organization.
@@
Figure 3: The annotation process of MEGA-Bench. We propose a “draft” taxonomy tree and then distribute the second-level nodes to annotators. We allow the annotators to gradually refine the tree structure as they add new tasks. Each task has many examples and a shared task-level instruction. Each example has a question and a ground truth answer.
@@
To ensure reliable commitment and annotation quality, we call up 16 designated expert annotators with rich LLM experience and computer science backgrounds. All annotators are graduate students or above with majors in computer science, electronics/communications engineering, bio-statistics, or finance, and 12 of them served as annotators or authors of LLM/VLM benchmarks published at top conferences. The annotators can 1) refine the “draft” taxonomy by adding/deleting nodes, 2) add “task group” nodes and then add a series of tasks under that, and 3) directly add tasks under an existing high-level node. We develop tools to facilitate the annotation process, including 1) an interactive annotation tool that defines the annotation format and automatically unifies all annotations into JSON files, 2) a GitHub repository to coordinate the task submission, reviewing, and discussion process, which was inspired by BIG-bench (Srivastava et al., 2022), and 3) a visualization tool that allows annotators to browse the existing tasks and the evaluation results of representative vision language models (VLMs). We coordinate all the annotators to ensure they understand our expectations and continuously improve our tools. Please see §B for complete details of annotation protocols.
@@
Task annotation. The annotation process contains two rounds. The annotators submit tasks to the benchmark by creating pull requests (PR) to the main branch of our GitHub repository. In the first round of the annotation process, we ask the annotators to contribute 20 tasks following the principles below to ensure the quality of the task:
@@
$\bullet$ Data source and output format: Creative tasks with diverse data sources and output formats are encouraged. If the data was collected from existing datasets, we ask annotators to adapt the original annotation into more specific questions and design more diverse answer formats.
@@
Before the second round of annotations, the core contributors review all tasks in the taxonomy tree and investigate the biases in the task distribution. We then host another annotator session to propose new meta-tasks to balance the distribution and maximize the coverage. We then distribute the updated tree nodes to annotators and employ the same guidelines to finish the second-round annotations. After this round, each annotator contributes at least another 30 tasks.
Quality control and refinement. We leverage commercial VLMs to examine the task quality. Concretely, we gather the results of GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro and compute an average score on each task. Tasks with almost 1.0 scores often have trivial questions (based on manual inspection) and can hardly distinguish the ability of different models. We ask the corresponding annotators to investigate and augment those tasks. For tasks with almost zero scores, the task reviewers audit them carefully and remove them if the zero score comes from incorrect annotations or insufficient instruction contexts. Finally, the benchmark contains a total of 505 tasks with roughly 8,200 examples, which is large enough to minimize the sample variance within each high-level taxonomy node. Please refer to §4.3 for an analysis of the number of examples per task.
@@ 3.2 Metrics for answers in diverse output formats
To properly evaluate the tasks with different output formats, we develop a set of highly-customized evaluation metrics in parallel with the benchmark construction process (§3.1). Figure 4 shows several examples of the model outputs along with the task’s associated metrics. When new tasks are submitted to our GitHub repository, we implement any new metrics specified by the task authors. We use two types of metrics: rule-based metrics and LLM-assisted metrics. All metrics are normalized into [0, 1], with 1.0 being the full mark.
Rule-based metrics. When there is a unique answer under the question context or the correctness of the answer can be verified by rules (e.g., if the generated story/poetry meets the desired formats or if the generated code can pass test cases), we implement rule-based metrics for evaluation. To satisfy the needs of all tasks submitted by annotators, we end up with a suite of over 40 rule-based metrics. Robust string parsing is also implemented to extract the answer from the model’s response. We conduct a sanity check to ensure the correct implementation of rule-based metrics. Specifically, we create an “oracle” model that always returns the ground truth, then compute its score over all tasks evaluated by rule-based metrics. The sanity check is passed when the “oracle” model gets a full 1.0 score. See §D.4 for details.
@@
Figure 4: Representative examples of MEGA-Bench’s diverse output formats and customized metrics (input queries are omitted). The outputs are extracted from real responses of GPT-4o (0513). We implement robust parsing to extract the final answer from raw model responses.
We divide the tasks into two subsets based on the different evaluation processes. The Core Set is evaluated with rule-based metrics to make the evaluation fast and cost-free. The Open-Ended Set is evaluated with metrics that use an LLM-as-a-judge, where the evaluation pipeline calls a proprietary LLM over an API. Specifically, we use GPT-4o-0806 (OpenAI, 2024a) as the judge LLM while maintaining an extensible implementation for using other judge models. The Core and Open-Ended sets contain 440 and 65 tasks, respectively.
@@ 3.3 Multi-dimensional Keywords for fine-grained analysis
Existing multi-task multimodal benchmarks analyze models according to dimensions like the image type and academic discipline (Yue et al., 2024a), ability (Liu et al., 2023b), or meta-task (Ying et al., 2024). MEGA-Bench offers a broad and diverse range of coverage across all these dimensions, and extends even further beyond them. As explained in §3.1, the taxonomy tree divides the tasks into general application scenarios, the most manageable dimension for distributing the annotation efforts to different annotators. After we collected all tasks and finished the quality control process, we grouped all tasks based on four extra dimensions: input visual type, input visual number, output format, and required skills (Figure 2). Each dimension has 6 to 10 keywords, enabling fine-grained analysis and comparison. Interactive visualization tools can then be developed based on our evaluation results, which allows model developers to delve deep into different aspects of a model and compare different models comprehensively.
@@
MathVista (Lu et al., 2023) Repurposed Diverse 1 Image MC / Num 1 5
MMBench (Liu et al., 2023b) Hybrid Diverse 1 Image Multi-choice (MC) 1 20
@@
Seed-Bench (Li et al., 2024c) New Existing Image/Video Multi-choice (MC) 1 12
VisIT-Bench (Bitton et al., 2023) Hybrid Diverse 1/2 Images Free-form (FF) 1 70
@@
MM-Vet (Yu et al., 2024b) Repurposed Existing 1 Image Free-form (FF) 1 16
MMMU (Yue et al., 2024a) New Diverse $\geq$1 Image MC / FF 1 30
@@
MileBench (Song et al., 2024) Repurposed Existing $>$1 Image MC / FF 2 12
@@
MMT-Bench (Ying et al., 2024) Repurposed Existing $\geq$1 Image/Video Multi-choice (MC) 1 162
MEGA-Bench New Diverse $\geq$1 Image/Video Unrestricted 45 505
@@ 3.4 Dataset statistics and comparison with other benchmarks
MEGA-Bench contains 505 real-world tasks with 8,186 manually annotated or repurposed samples. Even for repurposed data, considerable effort is needed to convert the original annotations into specific task descriptions, diverse output formats, and additional instructions to include auxiliary information about formatting. Figure 2 shows the task distribution of all five dimensions, and the detailed task taxonomy tree and statistics of each dimension are in Appendix C.
Table 1 compares MEGA-Bench to existing multimodal benchmarks. The key feature of our benchmark is the diversity across all aspects, driven by our high-level designs of diverse task applications and output formats. MMMU (Yue et al., 2024a; b) focuses on college-level exam questions with various discipline and image formats. All questions are single-image and answered in multiple-choice format. MMT-Bench (Ying et al., 2024) covers 162 concrete sub-tasks, enabling in-depth analysis based on their “taskonomy” and diverse input forms. However, all of the tasks MMT-Bench are from existing datasets, mostly under the “Perception” sub-tree in our taxonomy, and all outputs are in multiple-choice form like MMMU. To maximize task coverage and the diversity in model outputs with cost-effective inference, MEGA-Bench includes a much broader range of task types and output formats, while having fewer total samples compared to existing benchmarks.